### v1

# Risk Twin: Real-time Risk Visualization and Control for Structural Systems

Abstract
Digital twinning in structural engineering is a rapidly evolving technology that aims to eliminate the gap between physical systems and their digital models through real-time sensing, visualization, and control techniques. Although digital twins can offer dynamic insights into physical systems, their accuracy is inevitably compromised by uncertainties in sensing, modeling, simulation, and controlling. This paper proposes a specialized digital twin formulation, named Risk Twin, designed for real-time risk visualization and risk-informed control of structural systems. Integrating structural reliability and Bayesian inference methods with digital twining techniques, Risk Twin can analyze and visualize the reliability indices for structural components in real-time. To facilitate real-time inference and reliability updating, a “simulation-free” scheme is proposed, leveraging precomputed quantities prepared during an offline phase for rapid inference in the online phase. Proof-of-concept numerical and real-world Risk Twins are constructed to showcase the proposed concepts.
keywords:
Bayesian inference, digital twin, digital shadow, structural health monitoring, structural reliability
[…]
1 Introduction
Digital twin technology involves creating a virtual replica of a physical system to simulate, predict, and understand the behavior of the modeled system, facilitating decision-making under evolving conditions [1, 2, 3, 4]. This technology yields an increasing number of applications, such as civil infrastructures [5, 6, 7], mechanical [8, 9] and battery [10, 11] systems, medical and health [12, 13, 14], signal processing [15], manufacturing [16, 17], and machine learning [15, 18]. In structure and infrastructural systems, a highly relevant technology is structural health monitoring [19, 20], which focuses on identifying structural damage and deterioration over time. The boundary between digital twining and advanced structural health monitoring systems can be blurred, but in general, applying digital twining technology to structural engineering projects can yield interactive cyber-physical systems with broader applications than structural health monitoring. Ideally, the interactions between physical systems and their digital twins are bi-directional: sensor data collected from the physical system can lead to an updated digital model, and adjustments to control parameters in the digital model can also affect the physical system. For instance, Kapteyn et al. demonstrated the use of a digital twin to monitor and control the state of an unmanned aerial vehicle in real-time [21]. Torzoni et al. adapted the technique to civil structures, facilitating proactive decision-making regarding maintenance and management actions [22]. Furthermore, the interaction between physical and cyber components of a composite structure was successfully achieved by Xu et al. [23]. These studies highlight the dynamic interplay between digital and physical components, extending beyond simply representing the physical system in a digital format.
The inspiration for this study stems from the work of Haag and Anderl [24], where they demonstrated a pilot cyber-physical digital twin system of a bending beam using real-time sensing and three-dimensional computer-aided design visualization techniques. Their system accurately displayed the bending and deformation of the beam in real time, achieving a tight linkage between the physical and digital twins. However, a prevailing assumption in many digital twin systems is that the sensing, modeling, and simulation modules are deterministic [25, 26, 27, 28]. This assumption, however, is not realistic in real-world digital twin applications, where uncertainties are pervasive [29]. These uncertainties arise from various sources, including measurement uncertainties due to inaccuracies and variabilities of sensors; modeling uncertainties stemming from limitations and assumptions in computational models; and operational uncertainties, which involve variations in system performance under different real-world conditions and human interactions [30, 31]. For instance, stress-sensing devices in structural digital twin systems often exhibit high sensitivity to temperature and humidity, causing significant discrepancies in Analog-to-Digital (AD) conversion values for identical stress levels measured under differing conditions, such as noon versus night [32]. These uncertainties can compromise the accuracy of deterministic digital twin models. To properly quantify the influence of uncertainties in digital twin systems, Kapteyn et al. proposed a comprehensive framework based on Bayesian network [21]. This framework represents the states of physical systems and their digital counterparts as random variables and updates the latter through Bayesian network inference using the observed data. Moreover, two realistic applications of predictive digital twin are proposed to illustrate the framework, which have been further explored in subsequent research [12, 22].
Although the previously discussed methodologies can predict specific quantities of interest, there is an absence of a specialized framework for assessing the risk of structural safety on a unified scale. Such a framework is essential not only for quantifying uncertainties and risks within the structural system, but also for providing a holistic metric of risk across different structural components. For instance, a digital twin can monitor quantities such as stress, strain, and moment that collectively contribute to the safety of structural components, but it does not offer a directly comparable metric for the risk of each component. To fill this gap, this study introduces Risk Twin, a system designed to analyze and visualize the reliability indices for structural components in real-time, thereby facilitating risk-informed control.
Building on existing progress of integrating probabilistic analysis within the digital twinning technology, this study aims to further advance their application in risk assessment and control of civil structures. Specifically, this paper proposes (i) the concept of Risk Shadow as a digital reflection of structural risk; (ii) an efficient “simulation-free” method to update the Risk Shadow; and (iii) real-world benchmark Risk Twin systems. In the benchmark experiments, the bidirectional interaction between the physical and digital systems is demonstrated.
The paper is structured as follows: Section 2 introduces the basic concepts of Risk Twin. Section 4 presents implementation details for two benchmark Risk Twin systems. Additionally, demonstration videos for these Risk Twin systems are provided for further reference. Consequently, conclusive remarks are drawn in Section 5.
2 Risk Twin
If a digital twin operates with deterministic parameters, it will classify the state of a system as either “safe” or “failure.” However, uncertainties are prevalent in digital twin systems due to variabilities in sensing and modeling, hindering an exact reflection of the true state of a system. Risk twin is a probabilistic digital twin designed to provide real-time quantification, visualization, and control of risks for the modeled system. This section begins with a general introduction to the framework of Risk Twin and then delves into the detailed developments of its core components.
2.1 General framework
Risk Twin, illustrated in Figure 1, encompasses the forward information flow from the physical system to the digital model and the inverse flow from the digital model to the physical system. The forward flow includes (i) Bayesian inference – statistical inference for basic random variables, and (ii) Risk Shadow – the digital representation of risk. The inverse flow involves the risk-informed control of the physical system. The computational challenges lie in the efficient statistical inference for basic random variables and reliability indices. Specialized methods to address these challenges will be developed in subsequent sections.
Figure 1: Framework of the Risk Twin. The workflow begins with sensors collecting data from the physical twin. This data is then used to update the probabilistic model of basic random variables. The updated probabilistic model is propagated into a reliability analysis module to obtain reliability indices for system components. Subsequently, the Risk Shadow, displaying these reliability indices, is presented to users. Users can then decide whether to execute control operations upon perceiving the Risk Shadow. Once a control command is received, actuators, governed by an algorithm with operational cost constraints, act on the physical system, completing the full loop of information flows between the physical and digital systems. To address the computational challenge of real-time inference and reliability updating, the operation of risk twin is decomposed into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase.
2.2 Bayesian inference for the basic random variables
In a structural system, sources of randomness may include variations in material properties, geometric quantities, initial and boundary conditions, dynamic excitation, and environmental effects [33, 34, 31]. A computational model for the system introduces further uncertainty into these sources of randomness. Moreover, the control system, if present, may have additional random quantities. We define $\boldsymbol{X}\in\mathbb{R}^{n}$ to represent the basic random variables considered in the digital twin system. The term “basic” indicates that all the other random quantities of the modeled system are deterministic functions of $\boldsymbol{X}$. The distribution of $\boldsymbol{X}$ is denoted by $f(\boldsymbol{x})$. To ensure computational feasibility, we let the dimension of $\boldsymbol{X}$ be finite, implying that random fields and stochastic processes are represented by a finite number of basis functions. Let $\boldsymbol{Y}\in\mathbb{R}^{m}$ represent the measured quantities, and $\mathcal{M}:\boldsymbol{x}\in\mathbb{R}^{n}\mapsto\boldsymbol{y}_{\mathcal{M}}\in\mathbb{R}^{m}$ denote an end-to-end computational model that maps outcomes of $\boldsymbol{X}$ into predictions of $\boldsymbol{Y}$. The combined effect of modeling and measurement uncertainties is described by $f_{\boldsymbol{\epsilon}}(\boldsymbol{\epsilon})$, where $\boldsymbol{\epsilon}=\boldsymbol{y}-\boldsymbol{y}_{\mathcal{M}}$ is the difference between the measured $\boldsymbol{y}$ and the model prediction $\boldsymbol{y}_{\mathcal{M}}$. Note that to construct Bayesian inference for $\boldsymbol{X}$, the combined effect of modeling and measurement uncertainties should not be treated as a basic random variable. The posterior distribution of the basic random variables $\boldsymbol{X}$, given the knowledge of measurement $\{\boldsymbol{Y}=\boldsymbol{y}\}$, is:
$f(\boldsymbol{x}|\boldsymbol{y})\propto{f_{\boldsymbol{\epsilon}}(\boldsymbol{y}-\mathcal{M}(\boldsymbol{x}))}f(\boldsymbol{x})\,,$ (1)
where $f(\boldsymbol{x})$ and $f(\boldsymbol{x}|\boldsymbol{y})$ are respectively the prior and posterior distributions, and $f_{\boldsymbol{\epsilon}}(\boldsymbol{y}-\mathcal{M}(\boldsymbol{x}))$ is the likelihood function.
In practice, since data is collected in a temporal sequence, the posterior distribution may evolve with time. This process is expressed by the following recursive equation:
[…]
where subscripts such as $t$ and $1:t$ are introduced to explicitly show the evolving nature of the conditional distributions, $\boldsymbol{y}^{(1:t)}$ denotes data accumulated from a reference time point “1” to the current $t$, $f(\cdot|\boldsymbol{y}^{(1:t)})$ is the posterior to be established at the current step, and $f(\cdot|\boldsymbol{y}^{(1:t-1)})$ is the posterior from the previous step. The first two lines of Eq. (2) are simply marginalization and conditioning rules. The last line assumes a conditional independence structure similar to that in hidden Markov models.
The transition law for basic random variables, $f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})$ in Eq. (2), varies with the nature of $\boldsymbol{X}$. If the distribution of $\boldsymbol{X}$ remains constant over time, for instance, when $\boldsymbol{X}$ represents the elastic moduli or geometries of a structural system, we can assign $f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})=\delta(\boldsymbol{x}^{(t)}-\boldsymbol{x}^{(t-1)})$. Conversely, if the distribution changes significantly over time, such as when $\boldsymbol{X}$ represents the location and magnitude of occasional loads like wind and snow, we can assign $f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})=f(\boldsymbol{x}^{(t)})$. These two extreme scenarios for $f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})$ can simplify Eq. (2) into:
$f(\boldsymbol{x}^{(t)}|\boldsymbol{y}^{(1:t)})\propto\left\{\begin{aligned} &f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{x}^{(t)}))f(\boldsymbol{x}^{(t)}|\boldsymbol{y}^{(1:t-1)})\,,&&\text{if }f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})=\delta(\boldsymbol{x}^{(t)}-\boldsymbol{x}^{(t-1)})\\ &f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{x}^{(t)}))f(\boldsymbol{x}^{(t)})\,,&&\text{if }f(\boldsymbol{x}^{(t)}|\boldsymbol{x}^{(t-1)})=f(\boldsymbol{x}^{(t)})\end{aligned}\right.$ (3)
Introducing a hyperparameter $\alpha\in[0,1]$ to control the amount of information inherited from the previous posterior, Eq. (3) can be generalized into:
$f(\boldsymbol{x}^{(t)}|\boldsymbol{y}^{(1:t)})\propto f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{x}^{(t)}))\left(\alpha f(\boldsymbol{x}^{(t)}|\boldsymbol{y}^{(1:t-1)})+(1-\alpha)f(\boldsymbol{x}^{(t)})\right)\,.$ (4)
Depending on the specifications of $\boldsymbol{X}$ and $\boldsymbol{Y}$, the inference equation can predict latent states of a system. The normalizing constant for Eq. (4), denoted by $c^{(t)}$, can be expressed by:
$c^{(t)}=$ $\frac{\alpha^{t-1}}{\prod_{i=1}^{t-1}c^{(i)}}\mathbb{E}_{\boldsymbol{X}}\left[\prod_{i=1}^{t}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(i)}-\mathcal{M}(\boldsymbol{X}))\right]+(1-\alpha)\sum_{i=2}^{t-1}\frac{\alpha^{t-i}}{\prod_{j=1}^{t-i}c^{(t-j)}}\mathbb{E}_{\boldsymbol{X}}\left[\prod_{j=0}^{t-i}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t-j)}-\mathcal{M}(\boldsymbol{X}))\right]$ (5)
$+(1-\alpha)\mathbb{E}_{\boldsymbol{X}}\left[f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{X}))\right]\,,t>1\,.$
Similarly, a generalized moment of interest, $\mathbb{E}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}\left[q(\boldsymbol{X})\right]$, can be expressed by:
$\mathbb{E}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}\left[q(\boldsymbol{X})\right]=$ $\frac{1}{c^{(t)}}\left(\frac{\alpha^{t-1}}{\prod_{i=1}^{t-1}c^{(i)}}\mathbb{E}_{\boldsymbol{X}}\left[q(\boldsymbol{X})\prod_{i=1}^{t}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(i)}-\mathcal{M}(\boldsymbol{X}))\right]\right.$ (6)
$+(1-\alpha)\sum_{i=2}^{t-1}\frac{\alpha^{t-i}}{\prod_{j=1}^{t-i}c^{(t-j)}}\mathbb{E}_{\boldsymbol{X}}\left[q(\boldsymbol{X})\prod_{j=0}^{t-i}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t-j)}-\mathcal{M}(\boldsymbol{X}))\right]$
$\left.+(1-\alpha)\mathbb{E}_{\boldsymbol{X}}\left[q(\boldsymbol{X})f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{X}))\right]\right)\,,t>1\,.$
2.3 Risk shadow
A Risk Shadow visualizes the reliability indices of structural components. The reliability index provides a unified, comparable metric of risk across various structural components. For a generic component, the reliability index at time step $t$, denoted by $\beta^{(t)}$, is defined as follows:
$\beta^{(t)}\coloneqq-\Phi^{-1}(\mathbb{P}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}(\mathcal{R}))\,,$ (7)
where $\Phi^{-1}$ is the standard normal inverse cumulative distribution function and $\mathbb{P}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}(\mathcal{R})$ is the probability of a risk event $\mathcal{R}$ measured by $f(\boldsymbol{x}|\boldsymbol{y}^{(1:t)})$. Without loss of generality, the event $\mathcal{R}$ is defined as:
$\mathcal{R}\coloneqq\{\boldsymbol{x}\in\mathbb{R}^{n}:G(\boldsymbol{x})\leq 0\}\,,$ (8)
where the limit-state function $G(\boldsymbol{x})$ hinges on a computational model that maps $\boldsymbol{x}$ into performance variables to determine whether $\mathcal{R}$ occurs. A typical limit-state function is $G(\boldsymbol{x})=\text{Capacity}(\boldsymbol{x})-\text{Demand}(\boldsymbol{x})$, although more general forms may be required for specialized applications [35].
To compute $\beta^{(t)}$, we express $\mathbb{P}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}(\mathcal{R})$ as:
$\mathbb{P}_{\boldsymbol{X}|\boldsymbol{y}^{(1:t)}}(\mathcal{R})=$ $\frac{\Phi(-\beta^{(0)})}{c^{(t)}}\left(\frac{\alpha^{t-1}}{\prod_{i=1}^{t-1}c^{(i)}}\mathbb{E}_{\boldsymbol{X}|\mathcal{R}}\left[\prod_{i=1}^{t}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(i)}-\mathcal{M}(\boldsymbol{X}))\right]\right.$ (9)
$+(1-\alpha)\sum_{i=2}^{t-1}\frac{\alpha^{t-i}}{\prod_{j=1}^{t-i}c^{(t-j)}}\mathbb{E}_{\boldsymbol{X}|\mathcal{R}}\left[\prod_{j=0}^{t-i}f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t-j)}-\mathcal{M}(\boldsymbol{X}))\right]$
$\left.+(1-\alpha)\mathbb{E}_{\boldsymbol{X}|\mathcal{R}}\left[f_{\boldsymbol{\epsilon}}(\boldsymbol{y}^{(t)}-\mathcal{M}(\boldsymbol{X}))\right]\right)\,,t>1\,,$
where $\beta^{(0)}$ is the reliability index associated with the prior distribution of $\boldsymbol{X}$. Eq. (9), Eq. (5), and Eq. (6) are specifically designed to facilitate a “simulation-free” updating of the Risk Twin. The general idea is that in an offline phase, we can precompute some samples and quantities, so that in the online phase of Risk Twin updating, we only need to plug in observations $\boldsymbol{y}^{(1:t)}$ into simple algebraic equations, resulting in a negligible computational cost. Specifically, in the offline phase, we need to compute $\beta^{(0)}$ using a rare event simulation method, and then prepare the datasets:
$\mathcal{D}=\{(\boldsymbol{x}^{(i)},\mathcal{M}(\boldsymbol{x}^{(i)}))\}_{i=1}^{N}\,,\boldsymbol{x}^{(i)}\sim f(\boldsymbol{x})\,,$ (10)
$\mathcal{D}_{\mathcal{R}}=\{(\boldsymbol{x}^{(i)},\mathcal{M}(\boldsymbol{x}^{(i)}))\}_{i=1}^{N^{\prime}}\,,\boldsymbol{x}^{(i)}\sim f(\boldsymbol{x}|\mathcal{R})\,.$
Using these prepared datasets and $\beta^{(0)}$, the expectations in Eq. (9), Eq. (5), and Eq. (6) can be readily approximated when observations $\boldsymbol{y}^{(i)}$ are collected, thereby achieving a simulation-free approach in the online phase. The accuracy of this simulation-free approach depends on the number of precomputed samples and the accumulation of time steps. This latter issue can be mitigated by periodically setting the Risk Twin offline to update $\beta^{(0)}$, $\mathcal{D}$, and $\mathcal{D}_{\mathcal{R}}$ with respect to the recent $f(\boldsymbol{x}|\boldsymbol{y}^{(1:t)})$. It is worth mentioning that the time step in $\boldsymbol{y}^{(1:t)}$ defines a computational mesh that may differ from the real time points for collecting sensor data. For example, one can aggregate sensor data collected at multiple small consecutive time steps into a single $\boldsymbol{y}^{(i)}$ to improve computational efficiency.
2.4 Human-Risk shadow interaction
We envision risk control being conducted through Human-Risk Shadow interactions. The optimal action $a^{\star}$ at a specific time step depends on the Risk Shadow and the constraints of the control system. We formulate $a^{\star}$ as the solution to the following optimization problem:
$a^{\star}$ $=\mathop{\arg\min}_{a\in\Omega_{a}}\text{Cost}(a,\text{Risk Shadow})$ (11)
[…]
where $\boldsymbol{h}$ and $\boldsymbol{g}$ are equality and inequality constraints from the control system, such as geometry and power constraints, $\tilde{\boldsymbol{\beta}}(a)$ denotes the estimated reliability index at the target structural component after an action, $\boldsymbol{\beta}_{0}$ is the reliability index constraints for the systems.It should be noted that $\tilde{\boldsymbol{\beta}}(a)$ can be approximated based on the up-to-date information because the future data is still unavailable.
3 Proof-of-Concept Numerical Experiment
to be added…
4 Benchmark Physical Experiments
This section details the implementation of two benchmark experiments to showcase the capabilities of RT. These illustrative experiments employ materials and equipment that are easily accessible to engineers. The first experiment features a simply supported plate, illustrating the real-time Bayesian inference (Section 2.2) for quantifying the uncertainty in position and magnitude of an external force. The second experiment is a cantilever beam controlled by a mechanical arm, showcasing the risk shadow (Section 2.3) and human-risk shadow interactions (Section 2.4). Video demonstrations of these two experiments are provided below.
4.1 A simply supported plate
In this experiment, we constructed a $1\,\text{m}\times 1\,\text{m}\times 0.05\,\text{m}$ acrylic plate supported by four rubber cushions, each embedded with a stress sensor, as illustrated in Figure 2. The sensors, rubber pads, and acrylic plate were affixed using adhesive to prevent slipping. A vertical load was applied to the plate, with the computational goal of quantifying the uncertainty in its magnitude and position.
Figure 2: A simply supported plate: (a) the physical system, (b) a sketch illustration, (c) the rubber pad, and (d) coordinate system for the plate. The physical system consists of a acrylic plate, rubber pads, DuPont wires, sensors, and TTL-USB connectors. The computational platform is developed within the Python environment.
4.1.1 Computational model
Given a weight $w$ located at $(u_{0},v_{0})$ according to the coordinate system depicted in Figure 2(d), the reaction forces at the four supports are:
$f_{A}$ $=\frac{w(l-u_{0})v_{0}}{l^{2}}\,,$ (12)
$f_{B}$ $=\frac{wu_{0}V_{0}}{l^{2}}\,,$
$f_{C}$ $=\frac{w(l-u_{0})(l-v_{0})}{l^{2}}\,,$
$f_{D}$ $=\frac{wu_{0}(l-v_{0})}{l^{2}}\,,$
where $l=0.75$m is the effective side length of the square plate. Therefore, the basic random variables are $\boldsymbol{X}\equiv[W,U_{0},V_{0}]$, the measured quantities are $\boldsymbol{Y}\equiv[F_{A},F_{B},F_{C},F_{D}]$, and Eq. (12) defines the computational model $\mathcal{M}:\boldsymbol{x}\in\mathbb{R}^{3}\mapsto\boldsymbol{y}_{\mathcal{M}}\in\mathbb{R}^{4}$ that predicts the measured quantities given an outcome of the basic random variables.
4.1.2 Bayesian inference
Uniform priors are assumed for both the location and magnitude of the weight, with the location being anywhere over the plate and the magnitude ranging within $[0,10]$ kg. It was found that varying temperatures can lead to significant fluctuations in sensor readings, as illustrated in Figure 3. Additionally, temporal drift also affects measurement accuracy. To address these issues, sensor calibration was conducted using standard weights, and the standard deviation of the sensor measurements was estimated to be $0.1$ kg. The implementation details of the measurement system is illustrated in Figure 3 (c). Assuming a zero-mean Gaussian distribution for the measurement uncertainty, the likelihood function $f_{\boldsymbol{\epsilon}}(\boldsymbol{y}-\mathcal{M}(\boldsymbol{x}))$ is:
$f_{\boldsymbol{\epsilon}}(\boldsymbol{y}-\mathcal{M}(\boldsymbol{x}))=\exp{\left(-\sum_{i=1}^{4}\frac{(y_{i}-\mathcal{M}_{i}(\boldsymbol{x}))^{2}}{2\times 0.1^{2}}\right)}\,,$ (13)
where $\mathcal{M}_{i}$ is expressed by the $i$-th line of Eq. (12), and $y_{i}$, $i=1,2,3,4$, are measured forces at supports A, B, C, D, respectively.
To perform inference, we prepared $10^{5}$ samples during the offline phase, and the $\alpha$ in Eq. (4) is set to $0.5$. Notice that the effect of $\alpha$ is to dilute the previous posterior, used to construct the current prior; the inference results remain insensitive to $\alpha$ as long as $\alpha\neq 1$. Using the posterior mean as a point representation, Figure 4 illustrates an inference result. Subsequently, we shift the weight among five target positions, as illustrated in Figure 4(c), and conduct the inference in real-time. The accuracy of these results is documented in Table 1, and a video demonstrating the process is available at https://youtu.be/vVuwe4H075k.
Table 1: Accuracy of Bayesian inference for $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ and $P_{5}$.
Position $\Delta d$ $\Delta d/l$ $\Delta W$ $\Delta W/W$
$P_{1}$ 0.018 2.40% 0.01 0.25%
$P_{2}$ 0.052 6.93% 0.11 2.75%
$P_{3}$ 0.113 15.07% 0.32 8.00%
$P_{4}$ 0.087 11.60% 0.23 5.75%
$P_{5}$ 0.023 3.07% 0.74 18.5%
Figure 3: Stress sensors: (a) flexible thin film pressure sensor with a measuring range of $0-10$ Kg; (b) the force-AD (Analog-Digital) value relationship, measured during daytime and nighttime, where $T$ denotes temperature; (c) the implementation details of the measurement system.
Figure 4: Inference of the position and magnitude of the weight: (a) the physical system, (b) the inference results for magnitude and location, with two standard deviations of the means, and (c) route for testing real-time inference. For (c), the weight was moved along the route $P_{1}\mapsto P_{2}\mapsto\cdots\mapsto P_{5}$. where the video recording can be accessed via https://youtu.be/vVuwe4H075k.
4.2 A cantilever beam controlled by a mechanical arm
[…]
Figure 5: A cantilever beam controlled by a mechanical arm: (a) the physical system, (b) a sketch illustration, (c) the schemes of mechanical arm with physically dimensional sizes, and (d) definition of plane coordinates.
4.2.1 Computational model
The geometric and mechanical properties of the beam and the mechanical arm are summarized in Table 2. The basic random variables considered in this experiment include the weight of the load, $W$, the maximum allowable stress of the beam, $\mathrm{\Sigma}_{\max}$, the maximum allowable torque of the mechanical arm, $M_{\max}$, the inaccuracies in motor rotations, $\Delta\Theta_{i}$, where $i=1,2,3$, and the vertical control force of the mechanical arm, $F_{c}$. The measured quantity is the reaction force at the support, denoted by $R$. Therefore, the basic random variables are $\boldsymbol{X}\equiv[W,\mathrm{\Sigma}_{\max},M_{\max},\Delta\Theta_{1},\Delta\Theta_{2},\Delta\Theta_{3},F_{c}]$, and the measured quantity is $Y\equiv R$. The model $\mathcal{M}$ that predicts the measured quantity given an outcome of the basic random variables is:
$\mathcal{M}(\boldsymbol{x})=\left\{\begin{aligned} &\frac{l}{l_{1}}w\,,&&\text{uncontrolled phase}\,,\\ &\frac{l}{l_{1}}w-\frac{l_{c}+l_{1}}{l_{1}}f_{c}\,,&&\text{controlled phase}\,,\end{aligned}\right.$ (14)
where $l_{c}$ is the horizontal location of the control force, expressed by
$l_{c}=u_{0}+\sum_{i=1}^{3}\left(a_{i}\sin(\sum_{j=1}^{i}(\theta_{j}+\Delta\theta_{j}))\right)\,,$ (15)
where $u_{0}$ is the horizontal position of the base of the mechanical arm, $a_{i}$ is the length of the $i$-th arm component, and $\theta_{i}$ is the rotational angle of the $i$-th motor subjected to uncertainty specified by $\Delta\Theta_{i}$. Clockwise rotation is considered as the positive direction and counterclockwise is the negative. Notice that $\theta_{i}$ is known from the control system of the mechanical arm.
Similar to the previous experiment, the likelihood function is:
$f_{\boldsymbol{\epsilon}}(y-\mathcal{M}(\boldsymbol{x}))=\exp{\left(-\frac{(y-\mathcal{M}(\boldsymbol{x}))^{2}}{2\times 0.1^{2}}\right)}\,.$ (16)
The limit-state function for the beam at any location $u\in(0,l_{2})$ is defined as:
$G_{b}(\boldsymbol{x};u)=\left\{\begin{aligned} &\sigma_{\max}-\frac{w(l_{2}-u)h_{b}}{2I}\,,&&\text{uncontrolled phase or $u>l_{c}$ controlled phase}\,,\\ &\sigma_{\max}-\frac{w(l_{2}-u)h_{b}}{2I}+\frac{f_{c}(l_{c}-u)h_{b}}{2I}\,,&&\text{$u<l_{c}$, controlled phase}\,,\end{aligned}\right.$ (17)
which indicates failure when the stress at any location $u$ exceeds $\sigma_{\max}$. The limit-state functions for each motor of the mechanical arm are defined as:
$G_{a}^{i}(\boldsymbol{x})=m_{\max}-m_{i}\,,$ (18)
where $m_{i}$ denotes the moment at the $i$-th motor, expressed by:
$m_{1}=\left\{\begin{aligned} &G_{m}\left(a_{1}\sin(\theta_{1}+\Delta\theta_{1})+\sum_{j=1}^{2}\left(a_{j}\sin(\sum_{k=1}^{j}(\theta_{k}+\Delta\theta_{k}))\right)\right)\equiv m_{G1}\,,&&\text{uncontrolled phase}\,,\\ &m_{G1}+f_{c}\sum_{j=1}^{3}\left(a_{j}\sin(\sum_{k=1}^{j}(\theta_{k}+\Delta\theta_{k}))\right)\,,&&\text{controlled phase}\,,\end{aligned}\right.$ (19)
$m_{2}=\left\{\begin{aligned} &G_{m}\left(a_{2}\sin(\sum_{j=1}^{2}(\theta_{j}+\Delta\theta_{j}))\right)\equiv m_{G2}\,,&&\text{uncontrolled phase}\,,\\ &m_{G2}+f_{c}\sum_{j=2}^{3}\left(a_{j}\sin(\sum_{k=1}^{j}(\theta_{k}+\Delta\theta_{k}))\right)\,,&&\text{controlled phase}\,,\end{aligned}\right.$ (20)
$m_{3}=\left\{\begin{aligned} &0\,,&&\text{uncontrolled phase}\,,\\ &f_{c}\left(a_{3}\sin(\sum_{j=1}^{3}(\theta_{k}+\Delta\theta_{k}))\right)\,,&&\text{controlled phase}\,,\end{aligned}\right.$ (21)
where $G_{m}$ is the weight of the motor.
Table 2: Parameters for the beam and mechanical arm.
Component Parameters Type Mean Standard Deviation
$l$ Deterministic 1.0 m -
$l_{1}$ Deterministic 0.4 m -
$l_{2}$ Deterministic 0.6 m -
$h_{b}$ Deterministic 0.002 m -
Beam $b_{b}$ Deterministic 0.04 m -
I Deterministic $2.67\times 10^{-11}\,\text{m}^{4}$
$E$ Deterministic 200GPa -
$\mathrm{\Sigma}_{\max}$ Lognormal 250Mpa 25MPa
$a_{0}$ Deterministic 0.052 m -
$a_{1}$ Deterministic 0.093 m -
$a_{2}$ Deterministic 0.093 m -
$a_{3}$ Deterministic 0.078 m -
Arm $G_{m}$ Deterministic 5.5 N -
$\Delta\Theta_{i}$, $i=1,2,3$ Normal $0^{\circ}$ $3^{\circ}$
$M_{\max}$ Lognormal $1.5\,N\cdot m$ $0.15\,N\cdot m$
$(u_{0},v_{0})$ Deterministic (0.3,-0.3) m -
4.2.2 Risk shadow
To construct the risk shadow, we prepared $10^{5}$ samples for $\mathcal{D}$ and $10^{3}$ samples for $\mathcal{D}_{\mathcal{R}}$ (recall Eq. (10)) during the offline phase, and the $\alpha$ in Eq. (4) is set to $0.5$. This sample size can be insufficient for highly reliable components. Therefore, if a probability is estimated to be zero through Eq. (9), we apply the first-moment second-order method to obtain a preliminary estimate for visualization. The first-moment second-order analysis has been adapted to incorporate the offline and online phases for rapid inference. Sensor data are transmitted to the computational platform at a rate of $10$ times per second, thus the risk shadow has an image refresh frequency of $10$ Hz. We visualize reliability indices in the risk shadow using the color map in Table 3. The color map was chosen as a preliminary illustration of the risk shadow; it may require further investigation regarding scalability and differentiability.
Snapshots of the risk shadow are illustrated by Figure 6. It is worth mentioning that the third motor does not bear any load when it operates in the free motion state, so we assigned a reliability index of “10” to suggest the upper bound for the risk shadow outputs. More demonstrations of this risk twin system can be found in videos at https://youtu.be/MBMhvgd8KKM and https://youtu.be/XeRB4-JCY8A.
Table 3: Color map for risk shadow.
Level of Risk Reliability index Color RGB
Safe $\beta\geq 3.7$ [Uncaptioned image] [124, 252, 0]
Low Risk $3.73.2\leq\beta<$ 3.7 [Uncaptioned image] [255, 255, 0]
Medium Risk $3.22.7\leq\beta<$ 3.2 [Uncaptioned image] [240, 150, 110]
High Risk $\beta<2.7$ [Uncaptioned image] [255, 0, 0]
Figure 6: Snapshots of the risk twin with: (a) physical system in the uncontrolled phase, (b) physical system in the controlled phase, (c) risk shadow of (a) and (d) risk shadow of (b). In (a), a 0.5 kg weight is placed on the right end, and the system is in the uncontrolled phase. In (b), a 0.2 kg weight is placed on the right end, while the system is under the control of the mechanical arm. In (c) and (d), the beam is divided into four segments, displaying the maximum reliability indices within each segment. Furthermore, the reliability indices of the three motors are also displayed.
4.2.3 Human-risk shadow interaction
To implement risk shadow-based control for this experiment, we propose cost-benefit control actions that satisfy geometrical and reliability constraints. Specifically, the geometrical constraint is
$u-u_{0}$ $=\sum_{i=1}^{3}\left(a_{i}\sin(\sum_{j=1}^{i}(\theta_{j}+\Delta\theta_{j}))\right)\,,$ (22)
$v-v_{0}$ $=\sum_{i=1}^{3}\left(a_{i}\cos(\sum_{j=1}^{i}(\theta_{j}+\Delta\theta_{j}))\right)\,,$
where $(u,v)$ is a generic target endpoint position of the mechanical arm and $\theta_{i}$ can be controlled.
The reliability constraint ensures that, after an increment $\delta\theta_{i}$ for $\theta_{i}$, the reliability index for each motor does not exceed a threshold $\beta_{0}$, i.e.,
$\beta(G_{a}^{i}(\boldsymbol{x};\delta\boldsymbol{\theta}))-\beta_{0}\geq 0,,i=1,2,3,$ (23)
where $G_{a}^{i}(\boldsymbol{x};\delta\boldsymbol{\theta})$ denotes the limit-state Eq. (18) with $\boldsymbol{\theta}$ replaced by $\boldsymbol{\theta}+\delta\boldsymbol{\theta}$.
For a small increment $\delta\boldsymbol{\theta}$, the expected energy cost can be approximated by:
$C(\delta\boldsymbol{\theta})\approx\mathbb{E}\left[\sum_{i=1}^{3}\lvert M_{i}\delta\theta_{i}\rvert\right]\approx\sum_{i=1}^{3}\lvert m_{i}(\Delta\theta=\boldsymbol{0})\delta\theta_{i}\rvert\,,$ (24)
where the expectation is evaluated using the first-order approximation by fixing the random variables $\Delta\boldsymbol{\Theta}$ to their mean values (zeros). Notice that the expressions for $m_{i}$, recall Eqs. (19)-(21), vary with the uncontrolled (free-motion) and controlled phases. Therefore, it is ideal to adopt a small $\delta\boldsymbol{\theta}$ for each control step. Specifically, the action to achieve a target endpoint position $(u_{c},v_{c})$ of the mechanical arm is discretized into a sequence of small increments $\delta\boldsymbol{\theta}^{(\tau)}$, such that at each step, the energy described by Eq. (24) is minimized while satisfying the constraints given by Eq. (22) and Eq. (23). The corresponding optimization problem is defined as:
$\delta\boldsymbol{\theta}^{(\tau)}$ $=\mathop{\arg\min}_{\delta\boldsymbol{\theta}}C(\delta\boldsymbol{\theta})\,,$ (25)
[…]
where the target locations $(u^{(\tau)},v^{(\tau)})$ for the step $\tau$ is set to:
$(u^{\tau},v^{(\tau)})=(u^{0},v^{(0)})+\tau\frac{(u_{c},v_{c})-(u^{(0)},v^{(0)})}{n_{\tau}}\,,\tau=1,2,...,n_{\tau}\,,$ (26)
where a linear interpolation between the target $(u_{c},v_{c})$ and initial $(u^{0},v^{(0)})$ locations is adopted. This simple linear interpolation, chosen for its simplicity, may not yield a globally minimized energy cost from $(u^{(0)},v^{(0)})$ to $(u_{c},v_{c})$. A shows the implementation details for controlling the endpoint of the mechanical arm to a target position.
Figure 7 displays the risk shadow of the mechanical arm in free motion, with more details illustrated by the demonstration video at https://youtu.be/XeRB4-JCY8A. Figure 8 illustrates a control process by human-risk shadow interaction, with more details at https://youtu.be/MsjypDyqM40.
Figure 7: Illustration of controlling the mechanical arm: (a) move the endpoint to the location $(0.20,-0.10)$; (b) risk shadow of (a); (c) move the endpoint to the location $(0.40,-0.20)$; (d) risk shadow of (c).
Figure 8: Human-risk shadow interaction: (a) control location $u_{c}=0.2$ m; (b) risk shadow of (a); (c) control location $u_{c}=0.3$ m; (d) risk shadow of (c); (e) control location $u_{c}=0.4$ m; (f) risk shadow of (e). To illustrate the interaction between human risk and shadow, we set a hypothetical goal to control the reliability indices of the beam so that they exceed $4.2$, while the reliability indices of the motors should be no less than $3.1$. The first attempt to control the beam at location $u_{c}=0.20$ m is shown in panels (a) and (b). As the mechanical arm was elevated, the reliability of the motor near the base decreased to $3.11$, indicating that continuing the control would be risky. Consequently, the user adjusted the control location to $u_{c}=0.30$ m and successfully completed the control task by further elevating the arm. The actual control location achieved was $u_{c}=0.31$ m, with an error of $1$ centimeter. Panels (e) and (f) illustrate another control attempt at $u_{c}=0.40$ m. In this scenario, the control action was mechanically more effective compared to the $u_{c}=0.30$ m scenario, but the reliability indices of the motors were lower.
5 Conclusions
This paper introduces and demonstrates the concept of Risk Twin, a digital twin system designed for real-time reliability visualization and reliability-informed control. The computational challenges of statistical inference and reliability updating are addressed by decomposing the service life of risk twins into offline and online phases. During the offline phase, computationally intensive models are simulated to prepare datasets and quantities for rapid, simulation-free inference in the online phase. To illustrate the proposed concept, two benchmark risk twin experiments are conducted using commonly available tools and materials. The first experiment demonstrates real-time Bayesian inference for quantifying the uncertainty in the position and magnitude of an external force on a simply supported plate. The second experiment showcases the risk shadow and human-risk shadow interactions using a cantilever beam controlled by a mechanical arm. Future research could explore the application and adaptation of risk twin technology to real-world civil structures and infrastructures, such as the real-time reliability-based control of wind turbines.
Acknowledgement
This work is partly supported by the New Faculties’ Basic Research Capability Enhancement Program through No.2-9-2022-010 at China University of Geosciences (Beijing). Dr. Zeyu Wang appreciates the instructive comments from Prof. Daniel Straub at Technical University of Munich(TUM), the assistance of Yi Ren at Qualcomm(Shanghai) for his support in sensor connecting and Zhuoqun Gao for her meticulous care. These supports are really appreciated. Any opinions, findings, and conclusions expressed in this paper are those of the authors, and do not necessarily reflect the views of the sponsors.
Appendix A Implementation details of the mechanical arm control
Algorithm 1 Reliability-informed control of the mechanical arm
1. a. Read parameters, $\theta_{i}$, from mechanical arm with current position;
— b. Compute the current position $\boldsymbol{p}=(u^{0},v^{0})$ of the endpoint of the mechanical arm;
— c. Define the number of discretization $n_{\tau}$;
— d. Define $\boldsymbol{p}_{c}=(u_{c},v_{c})$.
2. Compute $\Delta{\boldsymbol{p}}=(\boldsymbol{p}_{c}-\boldsymbol{p})/n_{\tau}$;
3. If the mechanical arm touches the beam:
— a. Yes, compute $m_{i}$ using the controlled phase equation;
— b. No, compute $m_{i}$ using the uncontrolled phase equation.
4. $\boldsymbol{p}\leftarrow\boldsymbol{p}+\Delta{\boldsymbol{p}}$ and estimate $\delta{\boldsymbol{\theta}}^{\star}$ based on Eq. (25);
5. Control the arm moving to $\boldsymbol{p}$ using $\delta{\boldsymbol{\theta}}^{\star}$;
6. If $\boldsymbol{p}$ reaches $\boldsymbol{p}_{c}$:
— a. Yes, stop;
— b. No, go back to step 3.